# From Instruction to Event: Sound-Triggered Mobile Manipulation

### v1

@@ 6.3 Ablation Study and Further Discussion
Effect of Input Modality for the Navigate Skill. We evaluate the impact of different visual modalities, including head RGB, head depth, and their corresponding encoders for the Navigate skill (see Table 3(a)). First, the model relying solely on head depth with a CNN encoder yields the lowest performance. This underperformance arises because depth observations lack semantic texture, making it difficult for the model to learn meaningful associations between visual cues in depth observations and the semantic information contained in audio observations. Second, models using head RGB achieve better results, with the trainable CNN encoder performing slightly better than the frozen DINOv2 due to its training process that helps the model to adapt to the indoor environment. Third, integrating both head RGB and head depth yields the best overall performance. This combination proves effective because the modalities are complementary. Head RGB provides semantic context to align with audio sources, while depth offers precise geometric distance information. Despite the performance of the dual-modality method, we select head RGB with the DINOv2 encoder as our default configuration to maintain an optimal balance between performance and training efficiency.
@@
Effect of Input Modality for the Pick Skill. We investigate the impact of depth modalities on the Pick skill in Table 3(b), comparing the efficacy of head depth and arm depth. Using head depth alone yields the lowest success rate of 57.20%, primarily because the target object often falls outside the head camera’s optimal viewing range during manipulation. In contrast, the model relying solely on arm depth achieves a higher success rate of 66.66%, as the arm depth can directly observe the target when aligned. We observe that integrating both depth sources results in the best performance (88.73%). The two views prove complementary: the arm depth provides precise object-to-gripper distance, while the head depth offers global context to assist the arm in returning to the resting pose.
Effect of Input Modality for the Open Door Skill. For the Open Door skill, we validate the impact of different depth inputs, specifically analyzing head depth and arm depth (see Table 3(c)). Models relying on a single depth source achieve comparable performance, with the head depth only model (Model II) slightly outperforming the arm depth only model (Model I). We attribute this margin to the fact that head depth provides a global view of the door’s structure, whereas the arm camera often suffers from initial misalignment with the door’s specific affordance (i.e., the handle). Crucially, combining both modalities (Model III) results in a dramatic performance boost to 62.16%. This improvement stems from the complementary nature of the inputs: the head depth offers the global context necessary to locate the door frame, which in turn allows the agent to leverage the arm depth for precisely perceiving the target affordance for accurate manipulation.
7 Conclusion
@@
In this work, we introduce sound-triggered mobile manipulation, a novel paradigm that shifts agents from passive instruction-following to active, sound-driven interaction. To support this paradigm, we develop Habitat-Echo, a data platform integrating audio rendering with physical interaction, and establish a hierarchical baseline comprising a high-level task planner and low-level policies. Extensive experiments validate that the proposed baseline enables agents to actively localize and manipulate sound-emitting objects. Notably, in challenging dual-source scenarios, the baseline demonstrates the capability to interact with the primary source amidst acoustic interference from a secondary signal, and successfully proceed to manipulate the subsequent target.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Qualitative Result
We present qualitative visualization results for the four manipulation skills in Figure 5. Figure 5 (a) and (b) illustrate the Pick skill. In the success case (a), the agent grasps the object and correctly returns the arm to the resting position, while in the failure case (b), the arm fails to retract to the standard pose despite holding the object. The Place skill is depicted in Figure 5 (c) and (d). In Figure 5 (c), the object is stably placed on the table, whereas in (d), the object is released with an unstable pose and falls off the surface of the small table. Figure 5 (e) and (f) visualize the Open Door skill. In the success scenario (e), the arm pulls the door open and resets to the resting position, but in the failure scenario Figure 5 (f), the arm remains extended after opening the door. Finally, Figure 5 (g) and (h) show the Close Sink skill, where the height of the blue water column visually correlates with the faucet handle angle. In Figure 5 (g), the faucet is successfully closed, while in Figure 5 (h), the arm accidentally collides with the handle, causing the opening angle to increase rather than decrease.
Figure 5: Additional Qualitative Results of Different Skills. (a), (b) Success and failure cases of the Pick skill. (c), (d) Success and failure cases of the Place skill. (e), (f) Success and failure cases of the Open Door skill. (g), (h) Success and failure cases of the Close Sink skill.
Appendix B Prompt for Task Planning
To ensure a fair comparison, we utilize a standardized prompt structure across all task planners. The input consists of two components: a system prompt and a user prompt. The system prompt defines the detailed rules and logic for task planning, while the user prompt provides the environmental context, including visual and acoustic observations.
Prompt for Task Planning in SonicStow and SonicInteract. The system prompt for SonicStow and SonicInteract is structured into four components. The first component defines the strictly allowed skill vocabulary. The second establishes the logical mapping between audio and the corresponding skill chains. The third specifies the planning constraints and the mandatory JSON output format. Finally, the fourth component provides a structural template for the response. Complementing this, the user prompt supplies the visual and acoustic observations to ground the planning process.
(I) Prompt for Task Planning in SonicStow and SonicInteract System Prompt: You are an expert Embodied AI Task Planner. Your goal is to generate a strictly ordered list of skills based on the input auditory (audio) and visual (head_rgb) observations. 1. ALLOWED SKILL VOCABULARY: You are strictly limited to ONLY the following skills. Do not invent new words or change spellings: [“nav”, “pick”, “place”, “open_door”, “close_sink”]. 2. LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound and classify the source object to determine the skill sequence: 1. Category A: SonicStow Objects (a) If the sound source is: Mechanical Alarm, Phone (ringing tone), or Furby (toy). REQUIRED PLAN: [“nav”, “pick”, “place”] 2. Category B: SonicInteract Objects (a) If the sound source is: Doorbell. REQUIRED PLAN: [“nav”, “open_door”] (b) If the sound source is: Water Running. REQUIRED PLAN: [“nav”, “close_sink”] 3. CONSTRAINTS and OUTPUT FORMAT: 1. Never return an empty list. 2. Output MUST be a valid JSON object containing a single key “plan”. 3. Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string. 4. OUTPUT FORMAT: {“plan”: [“skill_1”, “skill_2”, …]}
User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”
Prompt for Task Planning in Bi-Sonic Manipulation. For the Bi-Sonic Manipulation task, we organize the system prompt into five segments to handle the dual-source complexity. The initial segment lists the permissible skill vocabulary. The second details the correspondence between sound categories and their required skill sequences. Crucially, the third segment enforces interaction priority. We explicitly identify the primary sound source, requiring the planner to deduce the secondary source from the mixed audio and generate a sequential execution plan for both. The fourth and fifth segments impose strict output constraints and provide a valid JSON template to ensure parsing reliability. Finally, the user prompt provides the model with multimodal context, including visual frames and audio data.
(II) Prompt for Task Planning in Bi-Sonic Manipulation System Prompt: You are an expert Embodied AI Task Planner. Your goal is to generate a strictly ordered list of skills based on the input auditory observations. There are exactly two sound sources detected. 1. ALLOWED SKILL VOCABULARY (STRICT): You are strictly limited to ONLY the following skills: [“nav”, “pick”, “place”, “open_door”, “close_sink”]. Do not invent new words or change spellings. 2. LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound to determine the action sequence: 1. Category A: SonicStow Objects (a) Sources: “Mechanical_Alarm”, “Phone”, “Furby” $\rightarrow$ [“nav”, “pick”, “place”] 2. Category B: SonicInteract Objects (a) Source: “Doorbell” $\rightarrow$ REQUIRED PLAN: [“nav”, “open_door”] (b) Source: “Running-Water” $\rightarrow$ REQUIRED PLAN: [“nav”, “close_sink”] 3. OBJECT MAPPING INSTRUCTION (CRITICAL): You are provided with a partial hint about the audio content: One of the sound sources is CONFIRMED to be: ‘‘{obj_1}". You have to interact with it first. Your task is to identify the other sound source from the audio and generate the plan. You MUST map the objects exactly as follows: 1. Key “first_sound”: Generate the plan for ‘‘{obj_1}". 2. Key “second_sound”: Identify the SECOND sound source from the audio (the one that is NOT ‘‘{obj_1}") and generate its plan. 3. Do NOT output plans for any other potential objects unless they are detected as the second sound. 4. CONSTRAINTS and OUTPUT FORMAT: 1. Output MUST be a valid JSON object containing a single root key “plan”. 2. The keys inside “plan” MUST be exactly “first_sound” and “second_sound”. 3. Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string. 5. Expected JSON Structure: { “plan”: { “first_sound”: [“skill_1”, “skill_2”, …], # $\leftarrow$ Actions for the object assigned to ’first_sound’ “second_sound”: [“skill_1”, “skill_2”, …] # $\leftarrow$ Actions for the object assigned to ’second_sound’ } }
User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”

### v2

@@ 6.3. Ablation Study and Further Discussion
Effect of Visual Modality for Navigate Skill. We evaluate the impact of different visual modalities, including head RGB, head depth for the Navigate skill (see Table 3(b)). First, the model relying solely on head depth yields the lowest performance. This underperformance arises because depth observations lack semantic texture, making it difficult for the model to learn meaningful associations between visual cues in depth observations and the semantic information contained in audio observations. Second, models using Head RGB achieve better results. Third, integrating both Head RGB and Head depth yields the best overall performance. This combination proves effective because the modalities are complementary. RGB provides semantic context to align with audio sources, while depth offers precise geometric distance information. Despite the performance of the dual-modality method, processing depth requires an additional visual encoder, which introduces additional computational burden. Therefore, we select single-modality Head RGB as the default configuration to maintain an optimal balance between performance, training efficiency, and computational cost.
@@
Effect of Input Modality for Pick Skill and Open Door Skill. We investigate the impact of depth modalities on the Pick skill and Open Door Skill in Table 3(c), comparing the efficacy of head depth and arm depth. For the Pick skill, using head depth alone yields the lowest success rate of 57.20%, primarily because the target object often falls outside the head camera’s optimal viewing range during manipulation. In contrast, the model relying solely on arm depth achieves a higher success rate of 66.66%, as the arm depth can directly observe the target when aligned. We observe that integrating both depth sources results in the best performance (88.73%). A similar trend appears in the Open Door Skill, where combining both depth views reaches the optimal result of 62.16%, significantly outperforming either single-modality setup. The two views prove complementary: the arm depth provides precise object-to-gripper distance and local affordances, while the head depth offers a global context of the gripper and object.
7. Conclusion
@@
In this work, we introduce sound-triggered mobile manipulation, a novel paradigm that shifts agents from passive instruction-following to active, sound-driven interaction. To support this paradigm, we develop Habitat-Echo, a data platform integrating audio rendering with physical interaction, and establish a hierarchical baseline comprising a high-level task planner and low-level policies. Extensive experiments validate that the proposed baseline enables agents to actively localize and manipulate sound-emitting objects. Notably, in challenging dual-source scenarios, the baseline shows the capability to interact with the primary source amidst acoustic interference from a secondary signal, and successfully proceed to manipulate the subsequent target.